# CCaaLF: Concurrency Control as a Learnable Function

### v3

@@ 6.2. Stored Procedure Mode
CCaaLF and all the baselines support the stored procedure mode. We ran all the experiments five times and report the average results. Each algorithm runs for three hours to match the time Polyjuice requires to reach optimal performance. We use IC3 as the initial agent function for agent function optimization.
@@
Figure 7. Performance comparison between CCaaLF and baselines under the YCSB-extended and TPC-C workloads in the stored procedure mode.
Varying access pattern. We fix the contention level and system load while varying the access pattern using four different hotspot distributions. [^2] The values of 1 and 0 indicate hotspot and uniform access, respectively. The four access patterns, from the left to the right in Figure 7(a), are $(0,0,0,1,0,0,0,0,0,0)$, $(1,0,0,1,1,0,0,0,0,0)$, $(0,0,0,1,0,1,0,1,1,0)$, and $(1,1,0,0,1,0,0,1,1,0)$. Figure 7(a) shows the results for the YCSB-extended workload. The hotspot distributions, shown from left to right, shift from being skewed at the beginning of the transaction to being skewed at the end. We observe that CCaaLF has the best throughput, up to 3.32$\times$ higher than that of the second-best baseline (Figure 7(a) Pat. 1). The reason is that CCaaLF learns a set of optimized pipeline-wait actions, timeouts, and conflict detection actions, enabling efficient transaction pipelining for hotspots while reducing the overhead for non-conflicting operations.
@@
It can be seen that the performance gap between CCaaLF and the second-best baseline shrinks as the hotspot distribution shifts toward the end of the transaction. This is because the learned pipeline actions, which contribute most of our performance advantage in this setting, become less effective as the hotspots shift forward. As highlighted in existing works (DBLP:conf/sigmod/GuoWYY21/bamboo, ; DBLP:journals/pvldb/FaleiroAH17/EWV, ), the advantage of dirty reads over clean reads relies on the trade-off between the gain from reducing block time and the overheads from cascading aborts and other issues caused by dirty reads. Therefore, with the hotspots shifting forward, the gain becomes smaller as the transaction is near completion. In other words, the benefit of learned pipeline-wait actions becomes less significant.
Varying contention level. We study the impact of contention levels by fixing the thread number to 16 and setting the number of warehouses to one for high contention (Figure 7(b)) and four for low contention (Figure 7(c)) under the TPC-C workload. We use shaded areas to represent the interquartile range, i.e., the 25th to 75th percentiles of throughput across multiple runs, illustrating the fluctuations of throughput in different CC algorithms. We can see that CCaaLF achieves up to 1.14$\times$ throughput of the second-best baseline. Also, CCaaLF achieves a more stable learning process compared to Polyjuice. Note that Polyjuice relies on random genetic search, which leads to high fluctuations. In contrast, CCaaLF employs an online-trained surrogate model to guide optimization, achieving greater stability throughout the optimization process.
@@
Varying system load. Figure 7(d) shows the results with varying system loads. We fix the number of warehouses at one and increase the thread number from 1 to 16. There are two main observations. First, CCaaLF outperforms all the baselines under high system load (more than four threads). This is because CCaaLF optimizes a better combination of CC actions that allows for higher concurrency. Second, both CCaaLF and Polyjuice perform slightly worse (less than 6.2% drop in throughput) than Silo under low system load (one thread), despite converging to the same CC actions. We note that Polyjuice reported a similar performance drop. The reason is that the learning processes of CCaaLF and Polyjuice incur additional overheads, which do not exist in the non-learned Silo system.
(a) Contention level
(b) System load
Figure 8. Performance comparison between CCaaLF and baselines under the TPC-E workload in the stored procedure mode.
Large benchmark. We further use TPC-E to evaluate CCaaLF against a larger search space. For comparison purposes, we follow  (DBLP:conf/osdi/WangDWCW0021/polyjuice, ) and vary skewness by adjusting the Zipfian distribution of the updates to the SECURITY table. In Figure 8(a), we use 16 threads and vary the Zipfian factor $\theta$ from 0 to 6. As the Zipfian factor $\theta$ increases, accesses to the SECURITY table become more skewed, leading to more conflicts between transactions accessing this table. The results show that CCaaLF consistently outperforms both 2PL and IC3 in these settings. 2PL exhibits suboptimal performance due to the high conflict detection costs inherent in its lock maintenance, whereas IC3 suffers from performance degradation caused by unnecessary pipeline-wait actions derived from its full conflict graph. CCaaLF effectively mitigates both issues by reducing conflict detection costs through learned actions and optimizing the conflict graph to eliminate unnecessary waits.
In Figure 8(b), we fix the Zipfian factor to six and increase the number of threads from 1 to 16. A higher thread number increases the system load and hence the likelihood of transaction conflicts. We observe that while CCaaLF performs worse than Silo in low system load (thread number $\leq$ 8), it outperforms the baselines (up to $\times1$) in high system load (16 threads). As expected, the learning overhead of CCaaLF becomes more significant under low system loads, and becomes less noticeable when the overall performance is greatly improved due to CCaaLF’s ability in identifying efficient CC actions under a high system load.
@@
Figure 9. Analysis of CCaaLF in the stored procedure mode.
@@
Workload drift. We evaluate the throughput changes of CCaaLF in response to workload drifts under the TPC-C workload. We simulate workload drifts by changing the number of threads for different time periods, i.e., 1 thread for Setting 1 starting at minute 0, 16 threads for Setting 2 starting at minute 200, and 4 threads for Setting 3 starting at minute 400. Figure 9(a) compares the throughput changes between optimizing from the initial agent function (“Restart”) and optimizing from the previously learned function (“w/o Restart”). We can observe that CCaaLF with “w/o Restart” fails to achieve optimal performance within three hours in both Setting 2 and Setting 3. This is because our graph reduction search algorithm only involves cutting-edge and merging-node mutations. We note that in order to fully explore the conflict graph search space, the optimization must begin with the full conflict graph, containing all nodes and edges. Therefore, starting the optimization from a previously learned optimal agent function, where the conflict graph has already been reduced, restricts the exploration of search space, and leads to a suboptimal conflict graph for the new workload.
@@
Impact of threshold value. We use eight different TPC-C workload settings [^3] Denote $(x,y)$ as $x$ threads and $y$ warehouses, the eight settings are $(1,1)$, $(2,1)$, $(4,1)$, $(8,1)$, $(16,1)$, $(16,2)$, $(16,4)$, and $(16,8)$. and analyze how the threshold value affects the accuracy of drift detection when switching the workload among these settings. We use false positive rate and false negative rate as the accuracy metrics. For the false positive rate, we set the workload to one of the eight settings and report the percentage of experimental runs where a workload drift is falsely detected despite no actual workload drift occurring. For the false negative rate, we evaluate all $7\times 8$ possible workload drifts, starting from one of the eight settings and changing to another, and report the percentage of experimental runs where these drifts are not detected. As shown in Figure 9(b), the threshold value of 10% leads to robust detection (3.6% false negative rate) without over-triggering optimization processes (0% false positive rate).
@@
Impact of features. We conduct an ablation study to evaluate the impact of the two features in the optimized feature set for the TPC-C workload. These features are the transaction type and the number of executed SQLs. Figure 9(c) shows the results under two settings: one warehouse and four warehouses, with features progressively removed from the optimized feature set. It can be seen that both features are necessary for high performance.
@@
Impact of learned actions. We evaluate the impact of the learned actions by reverting the learned CC actions in the final learned agent function to the actions in the initial agent function (IC3). Figure 9(d) shows that different actions have different impacts. Among them, the $\mathcal{{F}}_{D}$ and $\mathcal{{F}}_{P}$ sub-functions (see Section 5.1) — ones that are prioritized in our optimization pipeline — account for the majority of the performance improvement (85.8% and 80.6% in the one warehouse and 4 warehouse settings, respectively). The reason is that, although dirty reads allow for a higher degree of concurrency between transactions, they also introduce the risk of dependency cycles that can severely degrade performance. By optimizing $\mathcal{{F}}_{D}$ and $\mathcal{{F}}_{P}$, CCaaLF mitigates this risk while ensuring high concurrency.
Optimization time. Figure 9(e) shows the cumulative distribution function of the optimization time for CCaaLF and Polyjuice to reach peak performance. We can observe that CCaaLF has a significantly lower average optimization time (9.67 minutes) compared to Polyjuice (1.78 hours), achieving an 11$\times$ speedup in optimization time. Further, the optimization time of CCaaLF exhibits a lower variance than that of Polyjuice. Unlike Polyjuice which adopts a random genetic search, CCaaLF adopts a graph reduction search guided by an online-trained surrogate model and hence incurs a shorter optimization time with lower variance.
@@ 6.3. Interactive Transaction Mode
We compare CCaaLF against three other baselines that support the interactive transaction mode, namely 2PL, Silo, and CormCC. In this mode, we use 2PL as the initial agent function for agent function optimization.
@@
Figure 10. Performance comparison between CCaaLF and baselines in the interactive transaction mode.
Varying read ratio and transaction length. To understand the impacts of the read ratio and transaction length, we perform two sets of experiments using the YCSB-extended workload. First, we fix the transaction length at 10 and vary the read ratio (i.e., the percentage of read operations) from 0% to 100%. Second, we fix the read ratio at 50% and vary the transaction length from 1 to 16. The results are presented in Figure 10(c) and Figure 10(d), respectively. The throughput of all algorithms increases as the read ratio increases or the transaction length decreases. This trend is expected as higher read ratios or fewer transaction operations result in fewer conflicts among transactions. Further, CCaaLF outperforms the second-best baselines by up to 1.7$\times$ in throughput when the read ratio is below 100%. While under read-only workloads, CCaaLF performs slightly worse (by 2.5%) than Silo. For the same reason as mentioned in Section 6.2, learning-based algorithms can learn complex actions that deliver high performance when there are many conflicts, but the learning overhead becomes significant when there are few conflicts.
(a) Impact of surrogate model
(b) Time to peak performance
Figure 11. Analysis of CCaaLF in the interactive transaction mode.
Impact of surrogate model. We evaluate the impact of the surrogate model on CCaaLF’s throughput by varying the model’s training time under the TPC-C workload. The results for one warehouse and four warehouses are shown in Figure 11(a). As expected, longer surrogate model training times improve the surrogate model’s accuracy. This, in turn, enables CCaaLF to achieve higher performance by identifying better agent functions.
Optimization time. Figure 11(b) depicts the cumulative distribution function of the optimization time for CCaaLF to reach peak performance. Compared to the stored procedure mode (Figure 9(e)), the optimization time in the interactive transaction mode is lower (3.59 minutes on average) because CCaaLF does not need to learn $\mathcal{{F}}_{P}$ for the fact that dirty reads are not permitted in this mode.
@@
Figure 12. Case studies for learned actions.
6.4. Case Study of Learned Actions
Figure 12 shows two examples that compare the CC actions learned by CCaaLF with those of the second-best baseline, i.e., Polyjuice in the stored procedure mode and 2PL in the interactive transaction mode.
@@ 6.4. Case Study of Learned Actions
In the stored procedure mode, as shown in Figure 12(a), both CCaaLF and Polyjuice read dirty data during WAREHOUSE table accesses and track the transaction dependency $T_{no}\to T_{pay}\to T^{\prime}_{no}$. Polyjuice learns to perform the STOCK table access of $T^{\prime}_{no}$ only after the STOCK table access of $T_{no}$, preventing a potential dependency cycle $T_{no}\to T_{pay}\to T^{\prime}_{no}\to T_{no}$. In contrast, CCaaLF learns a conflict graph where STOCK table accesses of new order transactions are unlikely to conflict with each other. Therefore, it enables a more efficient transaction interleaving where the STOCK table access from $T_{no}$, the DISTRICT table access from $T_{pay}$, and the STOCK table access from $T^{\prime}_{no}$ can be executed in parallel.
@@
In the interactive transaction mode (Figure 12(b)), 2PL resolves the conflicts between $T_{no}$ and $T_{pay}$ by performing the WAREHOUSE table access of $T_{no}$ only after $T_{pay}$ commits. This results in a long execution time. CCaaLF, on the other hand, learns to order $T_{no}$ before $T_{pay}$ ($T_{no}\to T_{pay}$), allowing $T_{no}$ to optimistically access the WAREHOUSE table without waiting for $T_{pay}$ to commit. It also learns to assign a high priority to the DISTRICT table access of $T_{no}$, ordering this access before that of $T_{pay}$. This prevents a dependency cycle $T_{no}\to T_{pay}\to T_{no}$. The safe interleaving of the two transactions allows them to commit faster in CCaaLF.
7. Related Works
@@
Adaptive CC algorithms are designed to achieve high performance across different workloads. They follow a classify-then-assign workflow: analyzing workload information, classifying transactions based on manually selected features, and assigning the most suitable CC algorithm for each type of transaction. The assignment step can be either heuristic-based (DBLP:conf/icde/Su0Z21/C3, ; DBLP:conf/mascots/TaiM96/AEP, ; DBLP:conf/icde/LinCZ18/ASOCC, ; DBLP:conf/usenix/TangE18/CormCC, ; DBLP:conf/sigmod/SuCDAX17/Tebaldi, ) or learning-based (DBLP:conf/cidr/TangJE17/ACC, ; DBLP:conf/osdi/WangDWCW0021/polyjuice, ).
@@
Among the algorithms adopting heuristic assignment, C3 (DBLP:conf/icde/Su0Z21/C3, ) uses locality-sensitive hashing (LSH) to cluster transactions with similar SQL texts, which tend to conflict on shared data objects. It applies 2PL for intra-group conflicts and OCC for inter-group conflicts. AEP (DBLP:conf/mascots/TaiM96/AEP, ) predicts whether a distributed transaction will encounter conflicts based on local data, then uses EWL when there are no conflicts, and uses PSL when there are. MCC (DBLP:conf/sosp/XieSLAK015/Callas, ) and Tebaldi (DBLP:conf/sigmod/SuCDAX17/Tebaldi, ) classify transactions based on prior workload knowledge, constructing an algorithm hierarchy that switches between RP, SSI, and 2PL, depending on conflict type. ASOCC (DBLP:conf/icde/LinCZ18/ASOCC, ) categorizes data into hot, warm, and cold based on the contention level, then uses OCC for cold data, 2PL for hot data, and early validation with retry for warm data. CCaaLF covers more CC designs than these algorithms, and it learns a fine-grained and flexible combination of them for better performance.
@@
CCaaLF has more similarities with the algorithms adopting learning-based assignment. Polyjuice (DBLP:conf/osdi/WangDWCW0021/polyjuice, ) classifies transaction operations by their access ID (a hard-coded unique transaction access identifier) and uses an evolutionary algorithm to optimize pipeline-wait actions and transaction back-offs. ACC (DBLP:conf/cidr/TangJE17/ACC, ) clusters transactions similar to PartCC (DBLP:journals/pvldb/KallmanKNPRZJMSZHA08/HStore, ) and assigns the best algorithm among OCC, 2PL, and PartCC to each cluster. For fast reconfiguration, it also trains a predictive model that predicts the optimal algorithm for each cluster, enabling automatic online reconfiguration with offline-trained models. CCaaLF supports more CC design choices than these algorithms, therefore it achieves higher performance across more diverse workloads. In particular, it does not require prior knowledge of the workloads for access ID assignment or transaction clustering.

### v4

@@ 7.2. Stored Procedure Mode
NeurCC and all the baselines support the stored procedure mode. We run all the experiments five times and report the average results. Each algorithm runs for three hours to match the time Polyjuice requires to reach optimal performance. We use IC3 as the initial function for function optimization.
@@
Figure 7. Performance comparison between NeurCC and baselines under the YCSB-extended and TPC-C workloads in the stored procedure mode.
[…]
@@
The performance gap between NeurCC and the second-best baseline decreases as the hotspot distribution shifts toward the end of the transaction. This is because the learned pipeline-wait actions, which contribute the most to the performance advantage of NeurCC, become less effective as the hotspots are closer to the end of the transactions. In particular, pipeline-wait actions allow dirty reads, but the advantage of dirty reads over clean reads depends on the trade-off between the gain from reducing block time and the overheads from cascading aborts and other issues (Guo et al., 2021; Faleiro et al., 2017). As the hotspots shifting towards the end, the gain from reducing block time becomes smaller.
[…]
@@
Varying system load. We fix the number of warehouses to one and vary the system load by increasing the number of threadsfrom 1 to 16. Figure 7(d) shows the results. There are three main observations. First, NeurCC outperforms all the baselines under high load (more than four threads). This is because our algorithm finds a better combination of CC actions that allows for higher concurrency. Second, both NeurCC and Polyjuice perform slightly worse (less than 6.2% drop in throughput) than Silo under low load (one thread), even though they converge to the same CC actions. This gap is also reported in Polyjuice (Wang et al., 2021). The reason is that the learning processes of NeurCC and Polyjuice incur additional overheads, which do not exist in the non-learned Silo system. Third, NeurCC achieves higher throughput than the baselines even when the workload does not saturate the system (e.g., with 4 threads). This is because NeurCC can still extract more concurrency from the workload than the baselines.
@@
(f) Impact of optimization stages
Figure 8. Analysis of NeurCC in the stored procedure mode.
[…]
@@
Workload drift. We simulate workload drifts under the TPC-C workload by changing the number of threads at different time intervals, i.e., 1 thread for Setting 1 (starting at minute 0), 16 threads for Setting 2 (starting at minute 200), and 4 threads for Setting 3 (starting at minute 400). Figure 8(a) compares the system throughput when optimizing from the initial function (“Restart”), versus when optimizing from the previously learned function (“w/o Restart”). It can be seen that NeurCC reacts quickly to workload drifts. Furthermore, the “w/o Restart” approach fails to achieve optimal performance within three hours in both Setting 2 and Setting 3. This is because our graph reduction search algorithm involves only edge removal and node merging mutations. We note that in order to fully explore the conflict graph search space, the optimization process must begin with the full conflict graph, containing all nodes and edges. Therefore, starting the optimization from a previously learned optimal function, where the conflict graph has already been reduced, restricts the exploration of search space, and consequently leads to a suboptimal conflict graph for the new workload.
@@
Impact of threshold value. We use eight different TPC-C workload settings [^4] Denote $(x,y)$ as $x$ threads and $y$ warehouses, the eight settings are $(1,1)$, $(2,1)$, $(4,1)$, $(8,1)$, $(16,1)$, $(16,2)$, $(16,4)$, and $(16,8)$. and evaluate how the threshold value affects the accuracy of drift detection when switching the workload among these settings. We use false positive rate and false negative rate as the accuracy metrics. For the false positive rate, we set the workload to one of the eight settings and report the percentage of the experiment runs where a workload drift is falsely detected despite no actual workload drift occurring. For the false negative rate, we evaluate all $7\times 8$ possible workload drifts, starting from one of the eight settings and changing to another, and report the percentage of the experiment runs where these drifts are not detected. As shown in Figure 8(b), the threshold value of 10% leads to robust detection (3.6% false negative rate) without over-triggering the optimization process (0% false positive rate).
@@
Impact of features. We conduct an ablation study to evaluate the impact of the two features in the optimized feature set for the TPC-C workload. These features are the transaction type and the number of executed SQLs. Figure 8(c) shows the results for settings with one and four warehouses, in which features are progressively removed from the optimized feature set. It can be seen that both features are necessary for high performance.
@@
Impact of learned actions. We evaluate the impact of the learned actions by reverting the learned CC actions in the final learned function to the actions in the initial function (IC3). Figure 8(d) shows that different actions have different impacts. Among them, the $\mathcal{{F}}_{D}$ and $\mathcal{{F}}_{P}$ sub-functions (see Section 5) — ones that are prioritized in our optimization pipeline — account for the majority of the performance improvement (85.8% and 80.6% in the one warehouse and 4 warehouse settings, respectively). The reason is that, although dirty reads allow for a higher degree of concurrency between transactions, they also introduce the risk of dependency cycles that can severely degrade performance. By optimizing $\mathcal{{F}}_{D}$ and $\mathcal{{F}}_{P}$, NeurCC mitigates this risk while ensuring high concurrency.
Impact of optimization stages. We evaluate how each optimization stage contributes to the agent function. Starting from the initial agent function, we measure the system throughput after each optimization stage. Figure 8(f) shows the results under TPC-C workloads with one and four warehouses. It can be seen that the highest improvements are achieved in the early stages. In particular, the largest performance gains are observed in Stage 1, where the optimization of $\mathcal{{F}}_{D}$ and $\mathcal{{F}}_{P}$ are prioritized. The later stages yield marginal improvements, indicating that the agent function has likely converged. The results demonstrate that the proposed optimization pipeline is effective: by prioritizing influential actions early, NeurCC achieves fast convergence, which is useful given the limited optimization time budgets.
[…]
@@ 7.3. Interactive Transaction Mode
We compare NeurCC against three other baselines that support the interactive transaction mode, namely 2PL, Silo, and CormCC. In this mode, we use 2PL as the initial function for function optimization.
@@
(e) Stress test
Figure 9. Performance comparison between NeurCC and baselines in the interactive transaction mode.
[…]
Rapid workload changes. Although real-world workloads typically evolve gradually (Ma et al., 2018; Das et al., 2016), rapid workload changes can still arise from planned events such as yearly promotional and sales events. To evaluate NeurCC’s robustness under such rapid changes, we vary the number of client threads from 1 to 16 at different switching frequencies. Figure 9(e) shows the average throughput over 1 hour. It can be seen that NeurCC performs consistently better than the baselines, even under highly unstable workloads. At the highest switching frequency, i.e. 1 second, NeurCC achieves 85.90 ktps, which is comparable to the best static baseline (Silo achieves 85.57 ktps). The results demonstrate that NeurCC is robust under rapid workload changes.
Figure 10. Time to peak performance in the interactive transaction mode.
[…]
Optimization time. Figure 10 depicts the cumulative distribution function of the optimization time for NeurCC to reach peak performance. Compared to the stored procedure mode (Figure 8(e)), the optimization time in the interactive transaction mode is lower (3.59 minutes on average) because NeurCC does not need to learn $\mathcal{{F}}_{P}$ for the fact that dirty reads are not permitted in this mode.
@@
(c) Low vs. high contention
Figure 11. Examples of the learned actions.
Case studies for learned actions.
7.4. Examples of Learned Actions
Figure 11(a) and Figure 11(b) show examples that compare the CC actions learned by NeurCC with those learned by the second-best baselines. While our implementation uses row-level locking, we show only table names in the examples for clarity.
@@ 7.4. Examples of Learned Actions
In the stored procedure mode, Polyjuice is the second-best baseline. As shown in Figure 11(a), both NeurCC and Polyjuice read dirty data during WAREHOUSE table accesses and track the transaction dependency $T_{no}\to T_{pay}\to T^{\prime}_{no}$. Polyjuice learns to perform the STOCK table access of $T^{\prime}_{no}$ only after the STOCK table access of $T_{no}$, preventing a potential dependency cycle $T_{no}\to T_{pay}\to T^{\prime}_{no}\to T_{no}$. In contrast, NeurCC learns a conflict graph where STOCK table accesses of $T_{no}$ and $T^{\prime}_{no}$ are unlikely to conflict with each other. Therefore, it enables a more efficient transaction interleaving where the STOCK table access from $T_{no}$, the DISTRICT table access from $T_{pay}$, and the STOCK table access from $T^{\prime}_{no}$ can be executed in parallel.
@@
In the interactive transaction mode (Figure 11(b)), 2PL is the second-best basline, and it resolves the conflicts between $T_{no}$ and $T_{pay}$ by performing the WAREHOUSE table access of $T_{no}$ only after $T_{pay}$ commits. This results in a long execution time. On the other hand, NeurCC learns to order $T_{no}$ before $T_{pay}$, allowing $T_{no}$ to optimistically access the WAREHOUSE table without waiting for $T_{pay}$ to commit. It also learns to assign a high priority to the DISTRICT table access of $T_{no}$, ordering this access before that of $T_{pay}$. This prevents a dependency cycle $T_{no}\to T_{pay}\to T_{no}$. The safe interleaving of the two transactions allows them to commit faster in NeurCC. This example demonstrates that by enabling better interleaving of transactions, NeurCC can outperform 2PL despite incurring extra validation costs.
Figure 11(c) compares the CC actions learned by NeurCC under two different contention levels. The system runs in interactive transaction mode, and we present the learned actions for both low contention (1 thread) and high contention (16 threads) settings. The optimal function varies significantly across workloads: under low contention, an OCC-like function performs best because of its low conflict detection overhead, whereas under high contention, a function that detects conflicts and resolves them immediately achieves superior performance.
8. Related Work
@@
Adaptive CC algorithms aim to achieve high performance across different workloads. They follow a classify-then-assign workflow: analyzing workload information, classifying transactions based on manually selected features, and assigning the most suitable CC algorithm for each type of transaction. The assignment step can be either heuristic (Su et al., 2021; Tai and Meyer, 1996; Lin et al., 2018; Tang and Elmore, 2018; Su et al., 2017; Huang et al., 2025; Hong et al., 2025) or based on learning (Tang et al., 2017; Wang et al., 2021; Zhuang et al., 2025).
@@
Among the algorithms adopting heuristic assignment, C3 (Su et al., 2021) uses locality-sensitive hashing (LSH) to cluster transactions with similar SQL texts, which tend to conflict on shared data objects. It applies 2PL for intra-group conflicts and OCC for inter-group conflicts. AEP (Tai and Meyer, 1996) predicts whether a distributed transaction will encounter conflicts based on local data, then uses EWL when there are no conflicts, and uses PSL when there are. MCC (Xie et al., 2015) and Tebaldi (Su et al., 2017) classify transactions based on prior workload knowledge, constructing an algorithm hierarchy that switches between RP, SSI, and 2PL, depending on conflict type. ASOCC (Lin et al., 2018) categorizes data into hot, warm, and cold based on the contention level, then uses OCC for cold data, 2PL for hot data, and early validation with retry for warm data. NeurCC covers more CC designs than these algorithms, and it learns a fine-grained and flexible combination of them for better performance.
@@
NeurCC shares similarities with other algorithms that adopt learning-based assignment. Polyjuice (Wang et al., 2021) classifies transaction operations by their access ID (a hard-coded unique transaction access identifier) and uses an evolutionary algorithm to optimize pipeline-wait actions and transaction back-offs. ACC (Tang et al., 2017) clusters transactions similar to PartCC (Kallman et al., 2008) and assigns the best algorithm among OCC, 2PL, and PartCC to each cluster. For fast reconfiguration, it also trains a model that predicts the optimal algorithm for each cluster, enabling automatic online reconfiguration with offline-trained models. NeurCC supports more CC design choices than these algorithms, and achieves better performance across more diverse workloads. Furthermore, it does not require prior knowledge of the workloads for access ID assignment or transaction clustering.